# Shortest-path percolation on scale-free networks

### v1

@@ Abstract
The shortest-path percolation (SPP) model aims at describing the consumption, and eventual exhaustion, of a network’s resources. Starting from a graph containing a macroscopic connected component, random pairs of nodes are sequentially selected, and, if the length of the shortest path connecting the node pairs is smaller than a tunable budget parameter, then all edges along such a path are removed from the network. As edges are progressively removed, the network eventually breaks into multiple microscopic components, undergoing a percolation-like transition. It is known that SPP transition on Erdős-Rényi (ER) graphs belongs to same universality class as of the ordinary bond percolation if the budget parameter is finite; for unbounded budget instead, the SPP transition becomes more abrupt than the ordinary percolation transition. By means of large-scale numerical simulations and finite-size scaling analysis, here we study the SPP transition on random scale-free networks (SFNs) characterized by power-law degree distributions. We find, in contrast with standard percolation, that the transition is identical to the one observed on ER graphs, denoting independence from the degree exponent. Still, we distinguish finite- and infinite-budget SPP universality classes. Our findings follow from the fact that the SPP process drastically homogenizes the heterogeneous structure of SFNs before the SPP transition takes place.
@@ I Introduction
Percolation theory offers comprehensive insights into networked complex systems by revealing the relationship between a system’s macroscopic connectedness and its microscopic structure. Traditionally, percolation theory has been studied to understand various physical phenomena such as diffusion process in porous media, expansion of forest fires, and gelation of molecules, providing profound understanding on the fundamental mechanisms underlying these systems in terms of phase transition and critical phenomena [1]. Later, percolation theory has been successfully adopted to characterize the robustness of various complex systems, represented as complex networks, against deletion or failure of microscopic elements, either nodes or edges [2, 3]. The key assumption linking percolation theory to robustness of complex systems is that global connectivity enables proper function.
The rule used to remove or add the microscopic elements of the network defines the specific percolation model. The ordinary percolation model, which removes microscopic elements uniformly at random, has been extensively studied [1]. While simple, this model has provided a theoretical foundation of understanding the organization of complex systems [4, 5]. Other percolation models and processes have been introduced and studied to consider more complex rules mimicking relevant scenarios in various complex systems. These models include targeted attacks [6], cascading failures on interdependent networks [7], explosive percolation [8], traffic percolation [9], triadic percolation [10], and extended-range percolation [11, 12].
@@
For example, the so-called “vanishing critical threshold problem”—where the non-percolating phase is reached only after removing every edge—has been observed when $\lambda<3$. This is related to the divergence of the second moment of the degree distribution $P(k)$ [5]. Another anomalous behavior is the breaking of the site-bond percolation universality [15]. Moreover, other studies claim contradictory results in the analytical expression of some critical exponents; notably the correlation exponent governing finite-size scaling (FSS) [5, 16, 15, 2], while recently reconciled in [17]. These anomalies are not only present in ordinary percolation models. For instance, the explosive percolation transition under the Achlioptas process turns out to display a wide range of different critical exponents, depending on whether the network’s degree distribution $P(k)$ follows a Poisson distribution or a power-law distribution with different values of $\lambda$ [8, 18, 19, 20, 21, 22]. These findings highlight the crucial role of network heterogeneity in the nature of the percolation transition.
@@
Recently, the shortest-path percolation (SPP) model, which is intended to describe the consumption and eventual depletion of network resources, was introduced and studied on ERNs [23]. In the model, one agent at a time demands a path from randomly selected origin and destination nodes; if the geodesic distance between the demanded pair of nodes does not exceed the budget $C$ available to the agent, then all edges along one of the shortest paths connecting origin and destination nodes are removed from the network; eventually, the network undergoes a percolation transition characterized by the disintegration of its macroscopic connected component into multiple microscopic clusters. $C$ is a tunable parameter of the SPP model; for $C=1$, the SPP reduces to the ordinary percolation model, however for $C>1$ the SPP model differs from the ordinary model as removed edges are topologically correlated. On ERNs, it was shown that the critical exponents of the SPP transition are different from the mean-field percolation exponents only if the parameter $C$ is unbounded [23, 24]; mean-field exponents of the ordinary percolation transition are found instead for the finite-$C$ case [23].
@@
In this paper, we study the SPP transition on SFNs by means of extensive Monte Carlo simulations and FSS analysis. For $C=1$, we clearly recover known results about the ordinary percolation model, where the transition on ERNs and SFNs are characterized by radically different behaviors as long as $\lambda<4$. For $C>1$ however, we find that the SPP critical exponents on SFNs are identical to those for ERNs; this happens regardless of the specific value of the degree exponent $\lambda$. We conjecture and confirm that it is due to the fact that the SPP process homogenizes the heterogeneous network structure before the SPP transition happens. Finally, we find that the correlation exponents vary from those for ERNs, depending on $\lambda$ and $C$, displaying a rich FSS behavior. We validate our numerical findings by performing data collapse across different system sizes.
@@ III.1 Uncorrelated Scale-free Networks
We apply the SPP model to random SFNs characterized by the power-law degree distribution, $P(k)\sim{k^{-\lambda}}$ if $k\in[k_{\min},k_{max}]$ and $P(k)=0$ otherwise, generated according to the so-called uncorrelated configuration model (UCM) [25]. We set the minimum degree $k_{\text{min}}=4$ to ensure that the resulting network is composed of a single connected component. To ensure negligible degree-degree correlations, we impose the structural cut-off $k_{\text{max}}=\sqrt{N}$ when $2<\lambda<3$ and the natural cut-off $k_{\text{max}}\sim N^{1/(\lambda-1)}$ for $\lambda\geq 3$.
@@ III.3 Finite-size scaling in the event-based ensemble
In our study, we rely on the so-called event-based ensemble [26, 27, 28]. Compared to the conventional ensemble which samples the critical observables at the same value of the control parameter $r=r_{c}$, the event-based ensemble samples critical observables at the single-instance pseudo-critical point $r=r_{c}^{*}(N)$ from each individual realization of the percolation process on a network of size $N$. The event-based ensemble is known to provide good estimates of critical exponents, especially when the percolation process entails substantial non-self-averaging effects [26, 27, 28].
Denote with $P(r,N)$ and $S(r,N)$ the values of the observables of Eqs. (1) and (2), respectively, when exactly a fraction $r$ of edges is removed from a network of size $N$ during the SPP process. The single-instance pseudo-critical point $r_{c}^{*}(N)$ is defined as
@@
The single-instance pseudo-critical point $r_{c}^{*}(N)$ is a random variable. Taking its ensemble average we get an estimate of the pseudo-critical point under the event-based ensemble
@@
where $r_{c}=r_{c}(\infty)$ is the critical point, $1/\bar{\nu}_{1}$ is the exponent related to the correlation length, and $b$ is a constant. The exponent $1/\bar{\nu}_{1}$ governs how fast the average $r_{c}(N)$ converges to its infinite-size limit $r_{c}$. However, in some cases the random variable $r_{c}^{*}(N)$ cannot be described in terms of the sole average: we also need information on the fluctuations
@@
where $\bar{\nu}_{2}$ is another critical exponent determining how fast the random variable $r_{c}^{*}(N)$ converges to its average. In general, $1/\bar{\nu}_{1}=1/\bar{\nu}_{2}$ for ordinary percolation on various lattice and network structures [26]. However, it was recently reported that $1/\bar{\nu}_{1}>1/\bar{\nu}_{2}$ for explosive percolation while $1/\bar{\nu}_{1}<1/\bar{\nu}_{2}$ for ordinary bond percolation on lattices above the upper-critical dimension with free-boundary conditions [29] and on SFNs [30].
The critical observables under the event-based ensemble can be obtained by taking the ensemble average as follows
@@
While the numerical estimate of $1/\bar{\nu}_{2}$ is readily obtained from Eq. (6), the estimation of the critical point $r_{c}$ and of the exponent $1/\bar{\nu}_{1}$ is a bit more elaborate. From the FSS ansatz in Eq. (5), the deviation of the pseudo-critical point $r_{c}(N)$ from the true critical point scales as $|r_{c}(N)-r_{c}|\sim N^{-1/\bar{\nu}_{1}}$. Taking the logarithm, this yields a linear relationship
@@ IV.1 Estimation of $\beta/\bar{\nu}$ and $\gamma/\bar{\nu}$
First, we estimate the critical exponent ratios $\beta/\bar{\nu}$ and $\gamma/\bar{\nu}$, which are respectively related to the relative size of the largest cluster $P$ and the average cluster size $S$ at criticality. In Fig. 1 (a)-(d), we plot $P_{c}(N)$ as a function of $N$ for different values of $\lambda$ and $C$. When $C=1$, the SPP model is identical to the ordinary bond percolation and we get $\beta/\bar{\nu}=0.54(1)$ and $\beta/\bar{\nu}=0.50(1)$ for $\lambda=2.1$ and $\lambda=2.7$, respectively. We can also obtain $\beta/\bar{\nu}=0.40(1)$ and $\beta/\bar{\nu}=0.35(1)$ for $\lambda=3.5$ and $\lambda=4.5$, respectively. These values nicely align with the theoretical prediction in  [5], which are $\lambda$-dependent. Surprisingly, regardless of the value of $\lambda$, we find that the estimated values are $\beta/\bar{\nu}\approx 0.33$ when $C=2$ and $C=3$, which is compatible with the mean-field value $\beta/\bar{\nu}=1/3$. Moreover, we find that for the infinite-$C$ SPP transition $\beta/\bar{\nu}\approx 0.22$, regardless of the value of $\lambda$. This value was reported in previous studies on ERNs [23]. It is also compatible with the analytical prediction $\beta/\bar{\nu}=0.2$ by Meng et al. [24].
@@
Figure 1: Estimation of $\beta/\bar{\nu}$ and $\gamma/\bar{\nu}$. We plot $P_{c}(N)$ as a function of $N$ for SFNs with (a) $\lambda=2.1$, (b) $\lambda=2.7$, (c) $\lambda=3.5$, and (d) $\lambda=4.5$. The slope of the straight lines at the log-log plot corresponds to the best estimate of the ratio $\beta/\bar{\nu}$. Similarly, we plot $S_{c}(N)$ as a function of $N$ for SFNs with (e) $\lambda=2.1$, (f) $\lambda=2.7$, (g) $\lambda=3.5$, and (h) $\lambda=4.5$. The slope of the straight lines at the log-log plot corresponds to the best estimate of the ratio $\gamma/\bar{\nu}$. Different symbols display results of the SPP model with $C=1$ (square), $C=2$ (circle), $C=3$ (triangle), and $C=N$ (diamond). Different line styles are used to display the best power-law fitting of the SPP model with $C=1$ (solid), $C=2$ (dashed), $C=3$ (dotted), and $C=N$ (dashed-dotted). Symbols are the average values and error bars are the standard deviations. We report the number of realizations in Table S1.
@@
Similar results are found for $\gamma/\bar{\nu}$ as well (see Fig. 1 (e)-(h)). In other words, our estimation reads $\gamma/\bar{\nu}\approx 0.33$ for $C=2$ and $C=3$, and $\gamma/\bar{\nu}\approx 0.55$ for infinite-$C$, regardless of $\lambda$. When $C=1$ and $\lambda<3$, $S_{c}(N)$ does not diverges as $N\to\infty$ due to the lack of the subcritical phase, thus displaying a decreasing behavior (see Fig. 1 (e) and (f)). When $\lambda=3.5$, we could obtain $\gamma/\bar{\nu}=0.16(1)$ and when $\lambda=4.5$, we could get $\gamma/\bar{\nu}=0.26(1)$. FSS theory predicts $\gamma/\bar{\nu}=1/5$ when $\lambda=3.5$ and $\gamma/\bar{\nu}=1/3$ when $\lambda=4.5$, respectively, but these discrepancies might originate from finite-size effects and sample-to-sample fluctuations from different network realizations. Note that the hyper scaling relation $2\beta/\bar{\nu}+\gamma/\bar{\nu}=1$ roughly holds for the estimated exponents across different $\lambda$ and $C$. We report details in Table 1.
@@
To verify our estimation, in Fig. 2, we show the data-collapse plots of the probability density function of rescaled critical observables using the estimated critical exponents $\beta/\bar{\nu}$ and $\gamma/\bar{\nu}$ for SFNs with $\lambda=2.7$ and various values of $C$. We could obtain a satisfactory data collapse across different system sizes, which implies that our estimations of $\beta/\bar{\nu}$ and $\gamma/\bar{\nu}$ are fair. We could obtain clear data-collapse plots for different values of $\lambda$ as well (see Figs. S1,  S2, and  S3).
@@
IV.2 Shortest-path percolation process homogenizes the network drastically

### v2

@@ Abstract
The shortest-path-percolation (SPP) model aims at describing the consumption and eventual exhaustion of a network’s resources. Starting from a network containing a macroscopic connected component, random pairs of nodes are sequentially selected, and if the length of the shortest path connecting the node pairs is smaller than a tunable budget parameter, then all edges along such a path are removed from the network. As edges are progressively removed, the network eventually breaks into multiple microscopic components, undergoing a percolation-like transition. It is known that the SPP transition on Erdős-Rényi networks (ERNs) belongs to the same universality class as of the ordinary bond percolation if the budget parameter is finite; for an unbounded budget, instead, the SPP transition becomes more abrupt than the ordinary percolation transition. By means of large-scale numerical simulations and finite-size scaling analysis, here we study the SPP transition on random scale-free networks (SFNs) characterized by power-law degree distributions. We find, in contrast with ordinary percolation, that the transition is identical to the one observed on ERNs, denoting independence from the degree exponent. Still, we distinguish finite- and infinite-budget SPP universality classes. Our findings follow from the fact that the SPP process drastically homogenizes the heterogeneous structure of SFNs before the SPP transition takes place.
@@ I Introduction
Percolation theory offers comprehensive insights into networked complex systems by revealing the relationship between a system’s macroscopic connectedness and its microscopic structure. Traditionally, percolation theory has been studied to understand various physical phenomena such as diffusion processes in porous media, the expansion of forest fires, and the gelation of molecules, providing a profound understanding on the fundamental mechanisms underlying these systems in terms of phase transition and critical phenomena [1]. Later, percolation theory has been successfully adopted to characterize the robustness of various complex systems, represented as complex networks, against deletion or failure of microscopic elements, either nodes or edges [2, 3]. The key assumption linking percolation theory to the robustness of complex systems is that global connectivity enables proper function.
The rule used to remove or add the microscopic elements of the network defines the specific percolation model. The ordinary percolation model, which removes microscopic elements uniformly at random, has been extensively studied [1]. While simple, this model has provided a theoretical foundation for understanding the organization of complex systems [4, 5]. Other percolation models and processes have been introduced and studied to consider more complex rules mimicking relevant scenarios in various complex systems. These models include targeted attacks [6], cascading failures on interdependent networks [7], explosive percolation [8], traffic percolation [9], triadic percolation [10], and extended-range percolation [11, 12].
@@
For example, the so-called “vanishing critical threshold problem”—where the nonpercolating phase is reached only after removing every edge—has been observed when $\lambda<3$. This is related to the divergence of the second moment of the degree distribution $P(k)$ [5]. Another anomalous behavior is the breaking of the site-bond percolation universality [15]. Moreover, other studies claim contradictory results in the analytical expression of some critical exponents; notably, the correlation exponent governing finite-size scaling (FSS) [5, 16, 15, 2], while recently reconciled in [17]. These anomalies are not only present in ordinary percolation models. For instance, the explosive percolation transition under the Achlioptas process displays a wide range of different critical exponents, depending on whether the network’s degree distribution $P(k)$ follows a Poisson distribution or a power-law distribution with different values of $\lambda$ [8, 18, 19, 20, 21, 22]. These findings highlight the crucial role of network heterogeneity in the nature of the percolation transition.
@@
Recently, the shortest-path-percolation (SPP) model, which is intended to describe the consumption and eventual depletion of network resources, was introduced and studied on ERNs [23]. In the model, one agent at a time demands a path from randomly selected origin and destination nodes; if the geodesic distance between the demanded pair of nodes does not exceed the budget $C$ available to the agent, then all edges along one of the shortest paths connecting origin and destination nodes are removed from the network; eventually, the network undergoes a percolation transition characterized by the disintegration of its macroscopic connected component into multiple microscopic clusters. $C$ is a tunable parameter of the SPP model; for $C=1$, the SPP reduces to the ordinary percolation model, however for $C>1$ the SPP model differs from the ordinary model as removed edges are topologically correlated. On ERNs, it was shown that the critical exponents of the SPP transition are different from the mean-field percolation exponents only if the parameter $C$ is unbounded [23, 24]; mean-field exponents of the ordinary percolation transition are found instead for the finite-$C$ case [23].
@@
In this paper, we study the SPP transition on SFNs by means of extensive Monte Carlo simulations and FSS analysis. For $C=1$, we clearly recover known results about the ordinary percolation model, where the transition on ERNs and SFNs is characterized by radically different behaviors as long as $\lambda<4$. For $C>1$ however, we find that the SPP critical exponents on SFNs are identical to those for ERNs; this happens regardless of the specific value of the degree exponent $\lambda$. We conjecture and confirm that it is because the SPP process homogenizes the heterogeneous network structure before the SPP transition happens. Finally, we find that the correlation exponents vary from those for ERNs, depending on $\lambda$ and $C$, displaying a rich FSS behavior. We validate our numerical findings by performing data collapse across different system sizes.
@@ III.1 Uncorrelated scale-free networks
We apply the SPP model to random SFNs characterized by the power-law degree distribution, $P(k)\sim{k^{-\lambda}}$ if $k\in[k_{\min},k_{max}]$ and $P(k)=0$ otherwise, generated according to the so-called uncorrelated configuration model (UCM) [25]. We set the minimum degree $k_{\text{min}}=4$ to ensure that the resulting network is composed of a single connected component. To ensure negligible degree-degree correlations, we impose the structural cutoff $k_{\text{max}}=\sqrt{N}$ when $2<\lambda<3$ and the natural cut-off $k_{\text{max}}\sim N^{1/(\lambda-1)}$ for $\lambda\geq 3$.
@@ III.3 Finite-size scaling in the event-based ensemble
In our study, we rely on the so-called event-based ensemble [26, 27, 28]. Compared to the conventional ensemble, which samples the critical observables at the same value of the control parameter $r=r_{c}$, the event-based ensemble samples critical observables at the single-instance pseudocritical point $r=r_{c}^{*}(N)$ from each individual realization of the percolation process on a network of size $N$. The event-based ensemble is known to provide good estimates of critical exponents, especially when the percolation process entails substantial non-self-averaging effects [26, 27, 28].
Denote with $P(r,N)$ and $S(r,N)$ the values of the observables of Eqs. (1) and (2), respectively, when exactly a fraction $r$ of edges is removed from a network of size $N$ during the SPP process. The single-instance pseudocritical point $r_{c}^{*}(N)$ is defined as
@@
The single-instance pseudocritical point $r_{c}^{*}(N)$ is a random variable. Taking its ensemble average, we get an estimate of the pseudocritical point under the event-based ensemble
@@
where $r_{c}=r_{c}(\infty)$ is the critical point, $1/\bar{\nu}_{1}$ is the exponent related to the correlation length, and $b$ is a constant. The exponent $1/\bar{\nu}_{1}$ governs how fast the average $r_{c}(N)$ converges to its infinite-size limit $r_{c}$. However, in some cases, the random variable $r_{c}^{*}(N)$ cannot be described in terms of the sole average: we also need information on the fluctuations
@@
where $\bar{\nu}_{2}$ is another critical exponent determining how fast the random variable $r_{c}^{*}(N)$ converges to its average. In general, $1/\bar{\nu}_{1}=1/\bar{\nu}_{2}$ for ordinary percolation on various lattice and network structures [26]. However, it was recently reported that $1/\bar{\nu}_{1}>1/\bar{\nu}_{2}$ for explosive percolation, while $1/\bar{\nu}_{1}<1/\bar{\nu}_{2}$ for ordinary bond percolation on lattices above the upper-critical dimension with free-boundary conditions [29] and on SFNs [30].
The critical observables under the event-based ensemble can be obtained by taking the ensemble average as follows:
@@
While the numerical estimate of $1/\bar{\nu}_{2}$ is readily obtained from Eq. (6), the estimation of the critical point $r_{c}$ and of the exponent $1/\bar{\nu}_{1}$ is a bit more elaborate. From the FSS ansatz in Eq. (5), the deviation of the pseudocritical point $r_{c}(N)$ from the true critical point scales as $|r_{c}(N)-r_{c}|\sim N^{-1/\bar{\nu}_{1}}$. Taking the logarithm, this yields a linear relationship:
@@ IV.1 Estimation of $\beta/\bar{\nu}$ and $\gamma/\bar{\nu}$
First, we estimate the critical exponent ratios $\beta/\bar{\nu}$ and $\gamma/\bar{\nu}$, which are respectively related to the relative size of the largest cluster $P$ and the average cluster size $S$ at criticality. In Figs. 1 (a)-(d), we plot $P_{c}(N)$ as a function of $N$ for different values of $\lambda$ and $C$. When $C=1$, the SPP model is identical to the ordinary bond percolation and we get $\beta/\bar{\nu}=0.54(1)$ and $\beta/\bar{\nu}=0.50(1)$ for $\lambda=2.1$ and $\lambda=2.7$, respectively. We can also obtain $\beta/\bar{\nu}=0.40(1)$ and $\beta/\bar{\nu}=0.35(1)$ for $\lambda=3.5$ and $\lambda=4.5$, respectively. These values nicely align with the theoretical prediction in  [5], which are $\lambda$-dependent. Surprisingly, regardless of the value of $\lambda$, we find that the estimated values are $\beta/\bar{\nu}\approx 0.33$ when $C=2$ and $C=3$, which is compatible with the mean-field value $\beta/\bar{\nu}=1/3$. Moreover, we find that for the infinite-$C$ SPP transition $\beta/\bar{\nu}\approx 0.22$, regardless of the value of $\lambda$. This value was reported in previous studies on ERNs [23]. It is also compatible with the analytical prediction $\beta/\bar{\nu}=0.2$ by Meng et al. [24].
@@
Figure 1: Estimation of $\beta/\bar{\nu}$ and $\gamma/\bar{\nu}$. We plot $P_{c}(N)$ as a function of $N$ for SFNs with (a) $\lambda=2.1$, (b) $\lambda=2.7$, (c) $\lambda=3.5$, and (d) $\lambda=4.5$. The slope of the straight lines on the log-log plot corresponds to the best estimate of the ratio $\beta/\bar{\nu}$. Similarly, we plot $S_{c}(N)$ as a function of $N$ for SFNs with (e) $\lambda=2.1$, (f) $\lambda=2.7$, (g) $\lambda=3.5$, and (h) $\lambda=4.5$. The slope of the straight lines on the log-log plot corresponds to the best estimate of the ratio $\gamma/\bar{\nu}$. Different symbols display results of the SPP model with $C=1$ (square), $C=2$ (circle), $C=3$ (triangle), and $C=N$ (diamond). Different line styles are used to display the best power-law fitting of the SPP model with $C=1$ (solid), $C=2$ (dashed), $C=3$ (dotted), and $C=N$ (dashed-dotted). Symbols are the average values and error bars are the standard deviations. We report the number of realizations in Table S1 of Ref. [31].
@@
Similar results are found for $\gamma/\bar{\nu}$ as well [see Figs. 1 (e)-(h)]. In other words, our estimation reads $\gamma/\bar{\nu}\approx 0.33$ for $C=2$ and $C=3$, and $\gamma/\bar{\nu}\approx 0.55$ for infinite $C$, regardless of $\lambda$. When $C=1$ and $\lambda<3$, $S_{c}(N)$ does not diverge as $N\to\infty$ due to the lack of the subcritical phase, thus displaying a decreasing behavior [see Figs. 1 (e) and (f)]. When $\lambda=3.5$, we could obtain $\gamma/\bar{\nu}=0.16(1)$ and when $\lambda=4.5$, we could get $\gamma/\bar{\nu}=0.26(1)$. FSS theory predicts $\gamma/\bar{\nu}=1/5$ when $\lambda=3.5$ and $\gamma/\bar{\nu}=1/3$ when $\lambda=4.5$, respectively, but these discrepancies might originate from finite-size effects and sample-to-sample fluctuations from different network realizations. Note that the hyperscaling relation $2\beta/\bar{\nu}+\gamma/\bar{\nu}=1$ roughly holds for the estimated exponents across different $\lambda$ and $C$. We report the details in Table 1.
@@
To verify our estimation, in Fig. 2, we show the data-collapse plots of the probability density function of rescaled critical observables using the estimated critical exponents $\beta/\bar{\nu}$ and $\gamma/\bar{\nu}$ for SFNs with $\lambda=2.7$ and various values of $C$. We could obtain a satisfactory data collapse across different system sizes, which implies that our estimations of $\beta/\bar{\nu}$ and $\gamma/\bar{\nu}$ are fair. We could obtain clear data-collapse plots for different values of $\lambda$ as well (see Figs. S1,  S2, and  S3 of Ref. [31]).
@@
IV.2 Shortest-path-percolation process homogenizes the network drastically